OpenStudy (anonymous):

F(x)=x^2+5 and h(x)=x+3 Find (f•h)(1/2)

OpenStudy (anonymous):

What is \(h(1/2)\)?

OpenStudy (anonymous):

Is it \[ (f\cdot h)\left(\frac 12\right) \]or\[ (f\circ h)\left(\frac 12\right) \]

OpenStudy (anonymous):

\[(f •h)(\frac{ 1 }{ 2 })\]

OpenStudy (anonymous):

What is \(•\) lol

OpenStudy (anonymous):

That's exactly how the question was written. I'm assuming multiply but I could not find the correct answer

OpenStudy (anonymous):

Do you know what \(h(1/2)\) is?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."